# Function-Correcting Codes for Insertion-Deletion Channels

### v1

@@ II-B Function-Correcting Codes(FCC)
[Irregular distance code] Let $D\in\mathbb{N}_{0}^{M\times M}$. A set of binary codewords $\mathcal{P}=\{p_{1},p_{2},\dots,p_{M}\}$ is called an irregular distance code if there exists an ordering of the codewords such that $d(p_{i},p_{j})\geq[I]_{ij}\quad\text{for all }i,j\in[M].$ We define $N(D)$ to be the smallest integer $r$ for which there exists a $D$-code of length $r$. If $I]_{ij}=D$ for all $i\neq j$, we write $N(M,D)$.
Lemma 5 ([10132545]).
@@
[Plotkin-like bound for irregular distance code] For any distance matrix $D\in\mathbb{N}_{0}^{M\times M}$,
$N(D)\geq\begin{cases}\dfrac{4}{M^{2}}\displaystyle\sum_{i,j;\,i<j}[I]_{ij},&\text{if $M$ is even},\\[11.99998pt] \dfrac{4}{M^{2}-1}\displaystyle\sum_{i,j;\,i<j}[I]_{ij},&\text{if $M$ is odd}.\end{cases}$
Lemma 6 ([Rajput2025]).
@@
For any locally $(\rho,\lambda)$-bounded function $f:\mathbb{F}_{2}^{k}\to S$, there exists a mapping $\mathrm{Col}_{f}:\mathbb{F}_{2}^{k}\to[\lambda]$ such that for any $u,v\in\mathbb{F}_{2}^{k}$ with $f(u)\neq f(v)$ and $d(u,v)\leq\rho$, we have $\mathrm{Col}_{f}(u)\neq\mathrm{Col}_{f}(v)$.
III Function-Correcting Codes for insertion-deletion errors
@@
In this section, we consider a function defined over the binary vector space $\mathbb{F}_{2}^{k}$, that is, $f:\mathbb{F}_{2}^{k}\rightarrow\operatorname{Img}(f)$, where the expressiveness of the function is given by $E=|\operatorname{Img}(f)|$. Let $x\in\mathbb{F}_{2}^{k}$ denote a binary vector on which the function $f$ is evaluated. The vector $x$ is transmitted over an asynchronous channel that may introduce at most $t$ errors, consisting of insertions, deletions, or both. To enable correction of such errors, $x$ is first encoded using an encoding function $\psi(x)=(x,p(x))$, where $p(x)\in\mathbb{F}_{2}^{r}$ represents the redundancy added to $x$ prior to transmission. Based on this encoding framework, we define function-correcting deletion codes, function-correcting insertion codes, and function-correcting insertion-deletion (insdel) codes, which are designed to recover the function value $f(x)$ in the presence of deletion, insertion, or combined insertion-deletion errors, respectively.
Definition 7 (Function-Correcting Deletion Codes(FCDC)).
@@
An encoding function $\psi(x):\mathbb{F}_{2}^{k}\rightarrow\mathbb{F}_{2}^{k+r}$ with $\psi(x)=(x,p(x))$, $x\in\mathbb{F}_{2}^{k}$ defines a function-correcting deletion codes for $f:\mathbb{F}_{2}^{k}\rightarrow\operatorname{Img}(f)$ if for all $x$ and $y$ such that $f(x)\neq f(y)$, following holds
$D_{t}(\psi(x))\cap D_{t}(\psi(y))=\phi$
@@
The above definition ensures that, for any two inputs $x$ and $y$ that produce distinct function values $f(x)$ and $f(y)$, their corresponding codewords remain distinguishable even after up to $t$ deletions. Consequently, any received subsequence resulting from at most $t$ deletions can be uniquely mapped back to a codeword, and hence to its corresponding function value $f(x)$. On the same line, function-correcting insertion codes can be defined as follows:
Definition 8 (Function-Correcting Insertion Codes(FCIC)).
@@
An encoding function $\psi(x):\mathbb{F}_{2}^{k}\rightarrow\mathbb{F}_{2}^{k+r}$ with $\psi(x)=(x,p(x))$, $x\in\mathbb{F}_{2}^{k}$ defines a function-correcting insertion codes for $f:\mathbb{F}_{2}^{k}\rightarrow\operatorname{Img}(f)$ if for all $x$ and $y$ such that $f(x)\neq f(y)$, following holds
$I_{t}(\psi(x))\cap I_{t}(\psi(y))=\phi$
Function-correcting insdel codes are defined with respect to the insdel metric as follows:
Definition 9 (Function-Correcting Insdel Codes(FCIDC)).
@@
The encoding map $\psi:\mathbb{F}_{2}^{k}\rightarrow\mathbb{F}_{2}^{k+r}$ defined as $\psi(x)=(x,p(x))$ yields a function-correcting code for the function $f:\mathbb{F}_{2}^{k}\rightarrow\text{Img}(f)$ in insdel metric if for all $x$ and $y$ such that $f(x)\neq f(y)$, the insdel distance $d_{ID}(\psi(x),\psi(y))>2t$.
@@
(1) $\Rightarrow$ (3). If $D_{t}(\psi(x))\cap D_{t}(\psi(y))=\varnothing$, then there is no common subsequence of length at least ntn-t for $\psi(x)$ and $\psi(y)$. Hence
@@
so $d_{\mathrm{ID}}(\psi(x),\psi(y))=2(n-\operatorname{LCS}(\psi(x),\psi(y)))>2t$. Hence every $t$-function-correcting deletion code is $t$-function-correcting insdel code.
(3) $\Rightarrow$ (1). If $d_{\mathrm{ID}}(\psi(x),\psi(y))>2t$, then $\operatorname{LCS}(\psi(x),\psi(y))<n-t$. Hence, there cannot exist a common subsequence of length $\geq n-t$. Therefore, we get
$(D_{t}(\psi(x))\cap D_{t}(\psi(y))=\varnothing.$
@@
this implies the existence of a common subsequence of length $\geq n-t$. This contradicts (1). Thus $I_{t}(\psi(x))\cap I_{t}(\psi(y))=\varnothing$.
@@
Given the equivalence among the three code formulations discussed above, we restrict our attention, in the remainder of this paper, to function-correcting insertion-deletion (insdel) codes.
Next, we define the optimal redundancy of a function-correcting insdel code (FCIDC) for a function $f$ as it is a key parameter and plays a pivotal role in the study of FCIDCs, as it fundamentally distinguishes them from the classically investigated error-correcting insdel codes by potentially achieving strictly lower redundancy for the same error tolerance.
Definition 10 (optimal redundancy).
@@
if $r$ is the smallest achievable redundancy length. The optimal redundancy is denoted by $r^{f}_{\mathrm{ID}}(k,t)$.
III-A Irregular Insdel-Distance Codes
In this section, we define irregular insdel-distance codes and establish their relationship with FCIDCs. Leveraging this connection, we derive several general results related to FCIDCs and further obtain both lower and upper bounds on their optimal redundancy.
We first define insdel distance matrices (Definition 11) for a function $f$, followed by irregular insdel-distance codes (Definition 12) in which the insdel-distance between each pair of codewords should satisfy individual distance constraints.
Definition 11 (Insdel Distance Matrices).
Let $M,t\in\mathbb{N}$. Consider $M$ binary vectors ${x_{1}},...,{x_{M}}\in{\mathbb{F}^{k}_{2}}$. Then, ${\textbf{I}_{f}^{(1)}}(t,{x_{1}},...,{x_{M}})$ and ${\textbf{I}_{f}^{(2)}}(t,x_{0},\ldots,x_{M-1})$ are $M\times M$ insdel distance matrices corresponding to function $f$ with entries as follows:
$\scriptstyle[\mathbf{I}_{f}^{(1)}(t,x_{1},\dots,x_{M})]_{ij}=\begin{cases}[2t+2-d_{\text{ID}}(x_{i},x_{j})]^{+},&\text{if }f(x_{i})\neq f(x_{j}),\\ 0,&\text{otherwise}.\end{cases}$
@@ III-A Irregular Insdel-Distance Codes
$\scriptstyle[\mathbf{I}_{f}^{(2)}(t,x_{1},\dots,x_{M})]_{ij}=\begin{cases}[2t+2+2k-d_{\text{ID}}(x_{i},x_{j})]^{+},&\text{if }f(x_{i})\neq f(x_{j}),\\ 0,&\text{otherwise}.\end{cases}$
@@
Definition 12 (Irregular insdel-distance codes).
@@
  • $\mathcal{P}$ is an irregular insdel distance code of type  $1$ for matrix $\mathbf{I}$ if there exists an ordering of $\mathcal{P}$ such that
    $d_{\mathrm{ID}}(p_{i},p_{j})\geq[\mathbf{I}]_{i,j}\quad\text{for all }1\leq i,j\leq M.$
  • $\mathcal{P}$ is an irregular insdel distance code of type $2$ corresponding to the matrix $\mathbf{I}$ if it is an irregular insdel distance code of type  $1$ and, in addition, its codeword length satisfies $r\geq K$.
@@
When $\mathbf{I}]_{i,j}=I$ for all $i\neq j$, where $I\in\mathbb{N}$, we write $N^{(1)}_{\mathrm{ID}}(M,I)$ and $N^{(2)}_{\mathrm{ID}}(M,I;K)$, respectively.
Example 3.
Let $p_{1}=(0),p_{2}=(1),p_{3}=(1)$ and $p_{4}=(0)$. Then, $\{p_{1},p_{2},p_{3},p_{4}\}$ is an irregular insdel distance code of type $1$ corresponding to matrix $\textbf{I}_{f}^{1}(t;x_{1},x_{2},x_{3},x_{4})$ of Example 2. Clearly, $N^{(1)}_{\mathrm{ID}}(\textbf{I}_{f}^{1}(t;x_{1},x_{2},x_{3},x_{4}))=1.$
Now, let $K=2$ then $p_{1}=(000),p_{2}=(111),p_{3}=(111)$ and $p_{4}=(000)$ is an irregular insdel distance code of type $2$ corresponding to matrix $\textbf{I}_{f}^{(2)}(t;x_{1},x_{2},x_{3},x_{4})$ of Example 2. In this case $N^{(2)}_{\mathrm{ID}}(\textbf{I}_{f}^{(2)}(t;x_{1},x_{2},x_{3},x_{4});2)=3.$
@@
for the left-hand side inequality, $LCS(x,y)$ tends to become larger than $LCS(x_{1},y_{1})+LCS(x_{2},y_{2})$ due to cross-part matches, i.e., symbols in $x_{1}$ matching symbols in $y_{2}$ or vice versa. However, these additional matches cannot exceed the length of the shorter string, i.e., they are bounded by $min(k,r)$. Hence,
$LCS(x,y)\leq LCS(x_{1},y_{1})+LCS(x_{2},y_{2})-min(k,r)$
Lemma 8.
@@
multiplying 2-2 to equation 7 we get:
$-2(LCS(x_{1},y_{1})+LCS(x_{2},y_{2})-min(k,r))$
$\leq-2\cdot LCS(x,y)$
@@
$d_{ID}(x_{1},y_{1})+d_{ID}(x_{2},y_{2})-2min(k,r)$
$\leq d_{ID}(x,y)$
$\leq d_{ID}(x_{1},y_{1})+d_{ID}(x_{2},y_{2})$
The following example shows that the lower bound given in Lemma 8 is tight.
Example 4.
Let $x=(0000,10111)\in\mathbb{F}_{2}^{9}$ and $y=(1111,00000)\in\mathbb{F}_{2}^{9}$. Then, $d_{ID}(0000,1111)=8,\quad d_{ID}(10111,00000)=8\text{ and }d_{ID}(000010111,111100000)=8$. Hence, $d_{ID}(000010111,111100000)=d_{ID}(0000,1111)+d_{ID}(10111,00000)-2\cdot 4$.
@@
where the redundancy $r$ is optimal, i.e., $r=r_{ID}^{f}(k,t)$. On the contrary, suppose that $N^{(1)}_{ID}(\bm{I}_{f}^{(1)}(t,x_{1},x_{2},...,x_{2^{k}}))>r_{ID}^{f}(k,t)$. This implies the existence of distinct indices $i,j\in\{1,\ldots,q^{k}\}$ such that $f({x_{i}})\neq f({x_{j}})$ and
@@
Consequently, from equation 1 we obtain
$d_{ID}(\psi({x_{i}}),\psi({x_{j}}))\leq d_{ID}({x_{i}},{x_{j}})+d_{ID}({p_{i}},{p_{j}})<2t+1.$
@@
Next, we establish the reverse inequality $r_{ID}^{f}(k,t)\leq N^{(2)}_{ID}(\bm{I}_{f}^{(2)}(t,x_{1},x_{2},...,x_{2^{k}});k)$. Let $\mathcal{P}=\{{p_{1}},\ldots,{p_{2^{k}}}\}$ be a $\bm{I}_{f}^{(2)}(t,x_{1},x_{2},...,x_{2^{k}})$ irregular insdel distance code of type 2 and length $N^{(2)}_{ID}(\bm{I}_{f}^{(2)}(t,x_{1},x_{2},...,x_{2^{k}});k)$, and define the encoding function
$\psi:\mathbb{F}_{q}^{k}\to\mathbb{F}_{q}^{k+r},\quad{x_{i}}\mapsto({x_{i}},{p_{i}}).$
For every pair $i,j\in\{1,\ldots,q^{k}\}$ with $f({x_{i}})\neq f({x_{j}})$, we have
@@
Example 5.
With the help of Example 3 and Theorem 1 we can bound the optimal redundancy of function $f$ defined in Example 2 as follows:
@@
III-B Simplified Redundancy Lower Bounds
Using a smaller set of information vectors one can obtain lower bound on the redundancy of FCIDCs as follows:
Corollary 1.
Consider a collection of $M$ distinct binary words $x_{1},x_{2},\ldots,x_{M}$ of length $k$. Then, for a function $f$, the optimal redundancy of an FCIC is lower bounded as follows:

### v2

@@ 2.2 Function-Correcting Codes(FCC)
[Irregular distance code] Let $\textbf{D}\in\mathbb{N}_{0}^{M\times M}$. A set of binary codewords $\mathcal{P}=\{p_{1},p_{2},\dots,p_{M}\}$ is called an irregular distance code if there exists an ordering of the codewords such that $d_{H}(p_{i},p_{j})\geq[\textbf{D}]_{ij}\quad\text{for all }i,j\in[M].$ We define $N(\textbf{D})$ to be the smallest integer $r$ for which there exists a D-code of length $r$. If $\textbf{D}]_{ij}=D$ for all $i\neq j$, we write $N(M,D)$.
Lemma 2.16 ([1]).
@@
[Plotkin-like bound for irregular distance code] For any distance matrix $\textbf{D}\in\mathbb{N}_{0}^{M\times M}$,
$N(\textbf{D})\geq\begin{cases}\dfrac{4}{M^{2}}\displaystyle\sum_{i,j;\,i<j}[\textbf{D}]_{ij},&\text{if $M$ is even},\\[11.99998pt] \dfrac{4}{M^{2}-1}\displaystyle\sum_{i,j;\,i<j}[\textbf{D}]_{ij},&\text{if $M$ is odd}.\end{cases}$
Lemma 2.17 ([16]).
@@
Let $f:\mathbb{F}_{2}^{k}\rightarrow\mathrm{Im}(f)$ be a locally $(\rho,\lambda)_{\mathrm{H}}$- function. Assume that $\mathrm{Im}(f)$ is equipped with a total order $\prec$, and that for every $u\in\mathbb{F}_{2}^{k}$, the set $B_{\mathrm{H}}^{f}(u,\rho)$ forms a contiguous block with respect to $\prec$. Then there exists a mapping
$\mathrm{Col}_{f}:\mathbb{F}_{2}^{k}\rightarrow[\lambda]$
such that for all $u,v\in\mathbb{F}_{2}^{k}$ satisfying $f(u)\neq f(v)\quad\text{and}\quad d_{H}(u,v)\leq\rho$,we have $\mathrm{Col}_{f}(u)\neq\mathrm{Col}_{f}(v)$.
3 Function-Correcting Codes for Insertion-Deletion Errors
@@
In this section, we consider a function defined over the binary vector space $\mathbb{F}_{2}^{k}$, that is, $f:\mathbb{F}_{2}^{k}\rightarrow{Im}(f)$, where the expressiveness of the function is given by E=|Im(f)|E=|{Im}(f)|. Let $x\in\mathbb{F}_{2}^{k}$ denote a binary vector on which the function $f$ is evaluated. The vector $x$ is transmitted over an asynchronous channel that may introduce at most $t$ errors, consisting of insertions, deletions, or both. To enable correction of such errors, $x$ is first encoded using an encoding function $\psi(x)=(x,p(x))$, where $p(x)\in\mathbb{F}_{2}^{r}$ represents the redundancy added to $x$ prior to transmission. Based on this encoding framework, we define function-correcting deletion codes, function-correcting insertion codes, and function-correcting insertion-deletion (insdel) codes, which are designed to recover the function value $f(x)$ in the presence of deletion, insertion, or combined insertion-deletion errors, respectively.
Definition 3.1 (Function-Correcting Deletion Codes(FCDCs)).
@@
An encoding function $\psi(x):\mathbb{F}_{2}^{k}\rightarrow\mathbb{F}_{2}^{k+r}$ with $\psi(x)=(x,p(x))$, $x\in\mathbb{F}_{2}^{k}$ defines a function-correcting deletion codes for $f:\mathbb{F}_{2}^{k}\rightarrow{Im}(f)$ if for all $x$ and $y$ such that $f(x)\neq f(y)$, following holds
$D_{t}(\psi(x))\cap D_{t}(\psi(y))=\emptyset$
@@
The aforementioned formulation guarantees that, even after up to $t$ deletions, the associated codewords for any two inputs $x$ and $y$ that result in different function values $f(x)$ and $f(y)$ remain recognizable. Consequently, any received subsequence resulting from at most $t$ deletions can be uniquely mapped back to a codeword, and hence to its corresponding function value $f(x)$. On the same line, function-correcting insertion codes can be defined as follows:
Definition 3.2 (Function-Correcting Insertion Codes(FCICs)).
@@
An encoding function $\psi(x):\mathbb{F}_{2}^{k}\rightarrow\mathbb{F}_{2}^{k+r}$ with $\psi(x)=(x,p(x))$, $x\in\mathbb{F}_{2}^{k}$ defines a function-correcting insertion codes for $f:\mathbb{F}_{2}^{k}\rightarrow\operatorname{Im}(f)$ if for all $x$ and $y$ such that $f(x)\neq f(y)$, following holds
$I_{t}(\psi(x))\cap I_{t}(\psi(y))=\emptyset$
Function-correcting insdel codes, designed to recover function evaluations in the presence of both insertion and deletion errors, are defined as follows.
Definition 3.3 (Function-Correcting Insdel Codes(FCIDCs)).
@@
The encoding map $\psi:\mathbb{F}_{2}^{k}\rightarrow\mathbb{F}_{2}^{k+r}$ defined as $\psi(x)=(x,p(x))$ yields a function-correcting code for the function $f:\mathbb{F}_{2}^{k}\rightarrow\text{Im}(f)$ in insdel metric if for all $x$ and $y$ such that $f(x)\neq f(y)$, the insdel distance $d_{ID}(\psi(x),\psi(y))>2t$.
@@
(1) $\Rightarrow$ (3). If $D_{t}(\psi(x))\cap D_{t}(\psi(y))=\emptyset$, then there is no common subsequence of length at least ntn-t for $\psi(x)$ and $\psi(y)$. Hence
@@
So $d_{ID}(\psi(x),\psi(y))=2(n-\operatorname{LCS}(\psi(x),\psi(y)))>2t$. Hence every $t$-function-correcting deletion code is $t$-function-correcting insdel code.
(3) $\Rightarrow$ (1). If $d_{ID}(\psi(x),\psi(y))>2t$, then $\operatorname{LCS}(\psi(x),\psi(y))<n-t$. Hence, there cannot exist a common subsequence of length $\geq n-t$. Therefore, we get
$(D_{t}(\psi(x))\cap D_{t}(\psi(y))=\emptyset.$
@@
This implies the existence of a common subsequence of length $\geq n-t$. This contradicts (1). Thus $I_{t}(\psi(x))\cap I_{t}(\psi(y))=\emptyset$.
@@
Given the equivalence among the three code formulations discussed above, we restrict our attention, in the rest of this paper, to function-correcting insertion-deletion codes (FCIDCs) .
Next, we define the optimal redundancy of an FCIDC for a function $f$ as it is a key parameter and plays a pivotal role in the study of FCIDCs.
Definition 3.5 (optimal redundancy).
@@
if $r$ is the smallest achievable redundancy length. The optimal redundancy is denoted by $r^{f}_{ID}(k,t)$.
3.1 Irregular Insdel-Distance Codes
In this section, we define irregular insdel-distance codes and establish their relationship with FCIDCs. Leveraging this connection, we derive several general results about FCIDCs and obtain both lower and upper bounds on their optimal redundancy.
We first define insdel distance matrices (Definition 3.6) for a function $f$, followed by irregular insdel-distance codes (Definition 3.10) in which the insdel-distance between each pair of codewords should satisfy individual distance constraints.
Definition 3.6 (Insdel Distance Matrices).
Let $M,t\in\mathbb{N}$. Consider $M$ binary vectors ${x_{1}},...,{x_{M}}\in{\mathbb{F}^{k}_{2}}$. Then, ${\textbf{I}_{f}^{(1)}}(t,{x_{1}},...,{x_{M}})$ and ${\textbf{I}_{f}^{(2)}}(t,x_{1},\ldots,x_{M})$ are $M\times M$ insdel distance matrices corresponding to function $f$ with entries as follows:
[…]
@@ 3.1 Irregular Insdel-Distance Codes
[…]
Remark 3.7.
Unlike the Hamming metric, the insdel metric allows cross-block alignment (Lemma 3.14) between concatenated words, leading to a loss of up to $2\min\{k,r\}$ in insdel distance. The additional $2k$ term in distance matrix $\bm{I}_{f}^{(2)}(t,x_{1},\dots,x_{M})]$ precisely compensates for this phenomena when $r\geq k.$
The following example illustrates that the +2k+2k term in the definition for $\bm{I}_{f}^{(2)}(t,x_{1},\dots,x_{M})$ is indispensible.
Example 3.8.
Let $k=4$ and $t=4$, and let $f:\mathbb{F}_{2}^{4}\to\mathrm{Im}(f)$ be a function such that $f(0000)\neq f(1111)$. Since
$d_{ID}(0000,1111)=2(4-0)=8,$
the distance requirement on the redundancy vectors corresponding to matrix $\mathbf{I}_{f}^{(1)}(t,x_{1},\dots,x_{M})$ reduces to
$d_{ID}\bigl(p(x_{1}),p(x_{2})\bigr)\;\geq\;2t+2-d_{ID}(x_{1},x_{2})=2.$
Choose the redundancy vectors
$p(x_{1})=101110,\qquad p(x_{2})=000000,$
for which
$d_{ID}\bigl(p(x_{1}),p(x_{2})\bigr)=8,$
and hence the $\mathbf{I}_{f}^{(1)}(t,x_{1},\dots,x_{M})$ condition is satisfied. Consider the concatenated codewords
$\psi(x_{1})=0000101110,\qquad\psi(x_{2})=1111000000.$
The longest common subsequence between these two words has length $6$, yielding
$d_{ID}\bigl(\psi(x_{1}),\psi(x_{2})\bigr)=2(10-6)=8<2t+2.$
Thus, despite satisfying the $\mathbf{I}_{f}^{(1)}(t,x_{1},\dots,x_{M})$ distance requirement, the resulting code fails to correct $t=4$ insertion–deletion errors.
In contrast, the distance requirement for the matrix $\mathbf{I}_{f}^{(2)}(t,x_{1},\dots,x_{M})$ enforces
$d_{ID}\bigl(p(x_{1}),p(x_{2})\bigr)\;\geq\;2t+2+2k-d_{ID}(x_{1},x_{2})=10,$
which excludes this choice of redundancy vectors. This example demonstrates that the additional $2k$ compensation term is essential to prevent distance collapse caused by cross LCS effects.
@@
Definition 3.10 (Irregular insdel-distance codes).
@@
  • $\mathcal{P}$ is an irregular insdel-distance code of type $1$ for matrix $\mathbf{I}$ if there exists an ordering of $\mathcal{P}$ such that
    $d_{ID}(p_{i},p_{j})\geq[\bm{I}]_{i,j}\quad\text{for all }1\leq i,j\leq M.$
  • $\mathcal{P}$ is an irregular insdel-distance code of type $2$ corresponding to the matrix $\mathbf{I}$ if it is an irregular insdel-distance code of type $1$ and, in addition, its codeword length satisfies $r\geq K$.
@@
When $\bm{I}]_{i,j}=I$ for all $i\neq j$, where $I\in\mathbb{N}$, we write $N^{(1)}_{ID}(M,I)$ and $N^{(2)}_{{ID}}(M,I;K)$, respectively.
Example 3.12.
Let $p_{1}=(0),p_{2}=(1),p_{3}=(1)$ and $p_{4}=(0)$. Then, $\{p_{1},p_{2},p_{3},p_{4}\}$ is an irregular insdel-distance code of type $1$ corresponding to matrix $\textbf{I}_{f}^{1}(t;x_{1},x_{2},x_{3},x_{4})$ of Example 3.9. Clearly, $N^{(1)}_{{ID}}(\textbf{I}_{f}^{1}(t;x_{1},x_{2},x_{3},x_{4}))=1.$
Now, let $K=2$ then $p_{1}=(000),p_{2}=(111),p_{3}=(111)$ and $p_{4}=(000)$ is an irregular insdel-distance code of type $2$ corresponding to matrix $\textbf{I}_{f}^{(2)}(t;x_{1},x_{2},x_{3},x_{4})$ of Example 3.9. In this case $N^{(2)}_{{ID}}(\textbf{I}_{f}^{(2)}(t;x_{1},x_{2},x_{3},x_{4});2)=3.$
@@
For the right-hand inequality, let $s$ be the longest common subsequence of $x$ and $y$. For each symbol in $s$ consider the positions to which it is matched in both $x$ and $y$. This induces a partition of the symbols of $s$ into four classes indexed by $(i,j),\quad i,j\in\{1,2\}$, where a symbol is of type $(i,j)$ if it is matched to $x_{i}$ in $x$ and to $y_{j}$ in $y$. Let $n_{ij}$ denote the number of symbols of type $(i,j)$ in $s$. Then following four cases can be considered:
Case 1: If the symbols are of type $(1,1)$ then they form a common subsequence of $x_{1}$ and $y_{1}$.Therefore,
$n_{11}\leq LCS(x_{1},y_{1}).$
Case 2: If the symbols are of type $(2,2)$ then they form a common subsequence of $x_{2}$ and $y_{2}$.Therefore,
$n_{22}\leq LCS(x_{2},y_{2}).$
Case 3: If the symbols are of type $(1,2)$ then they form a common subsequence of $x_{1}$ and $y_{2}$.Therefore,
$n_{12}\leq\min\{k,r\}$
Case 4: If the symbols are of type $(2,1)$ then they form a common subsequence of $x_{2}$ and $y_{1}$.Therefore,
$n_{21}\leq\min\{k,r\}$
We claim $n_{12}\cdot n_{21}=0,$ i.e., symbols of type $(1,2)$ and $(2,1)$ cannot appear together in $s$. On the contrary assume that there exists symbols $s_{i}$ and $s_{j}$ of type $(1,2)$ and $(2,1)$ of $s$, if $s_{i}$ preceeds $s_{j}$, then the ordering is respected in $x$ but violated in $y$. Similarly, if $s_{j}$ preceeds $s_{i}$, then the ordering is respected in $y$ but violated in $x$.
Hence, the claim follows and we have $n_{12}+n_{21}=\max\{n_{12},n_{21}\}\leq\min\{k,r\}.$
Lemma 3.14.
@@
multiplying (2)(-2) to equation 3.1 we get:
$-2(LCS(x_{1},y_{1})+LCS(x_{2},y_{2})+\min\{k,r\}$ $\leq-2\cdot LCS(x,y)$
@@
$d_{ID}(x_{1},y_{1})+d_{ID}(x_{2},y_{2})-2\cdot\min\{k,r\}\leq d_{ID}(x,y)\leq d_{ID}(x_{1},y_{1})+d_{ID}(x_{2},y_{2})$
Remark 3.15.
Lemma 3.14 suggests that cross-LCS effects cause the insdel distance between concatenated codewords to suffer a worst-case loss of $2\min\{k,r\}$ for redundancy length $r$. The resulting distance condition simplifies to demanding that the message vectors themselves already satisfy the full insdel distance constraint if $r<k$. This loss completely offsets any redundancy distance gain. Determining the type $2$ irregular insdel-distance code is motivated by the fact that nontrivial systematic function-correcting insertion-deletion codes only exist in the regime $r\geq k$.
The following example shows that the lower bound given in Lemma 3.14 is tight.
Example 3.16.
Let $x=(0000,10111)\in\mathbb{F}_{2}^{9}$ and $y=(1111,00000)\in\mathbb{F}_{2}^{9}$. Then,
$d_{ID}(0000,1111)=8,\quad d_{ID}(10111,00000)=8\text{ and }d_{ID}(000010111,111100000)=8$.
Hence, $d_{ID}(000010111,111100000)=d_{ID}(0000,1111)+d_{ID}(10111,00000)-2\cdot 4$.
@@
where the redundancy $r$ is optimal, i.e., $r=r_{ID}^{f}(k,t)$. On the contrary, suppose that
$N^{(1)}_{ID}(\bm{I}_{f}^{(1)}(t,x_{1},x_{2},...,x_{2^{k}}))>r_{ID}^{f}(k,t)$. This implies the existence of distinct indices $i,j\in\{1,\ldots,2^{k}\}$ such that $f({x_{i}})\neq f({x_{j}})$ and
@@
Consequently, from equation 3.2 we obtain
$d_{ID}(\psi({x_{i}}),\psi({x_{j}}))\leq d_{ID}({x_{i}},{x_{j}})+d_{ID}({p_{i}},{p_{j}})<2t+2.$
@@
Next, we establish the reverse inequality $r_{ID}^{f}(k,t)\leq N^{(2)}_{ID}(\bm{I}_{f}^{(2)}(t,x_{1},x_{2},...,x_{2^{k}});k)$. Let $\mathcal{P}=\{{p_{1}},\ldots,{p_{2^{k}}}\}$ be a $\bm{I}_{f}^{(2)}(t,x_{1},x_{2},...,x_{2^{k}})$ irregular insdel-distance code of type 2 and length $N^{(2)}_{ID}(\bm{I}_{f}^{(2)}(t,x_{1},x_{2},...,x_{2^{k}});k)$, and define the encoding function
$\psi:\mathbb{F}_{2}^{k}\to\mathbb{F}_{2}^{k+r},\quad{x_{i}}\mapsto({x_{i}},{p_{i}}).$
For every pair $i,j\in\{1,\ldots,2^{k}\}$ with $f({x_{i}})\neq f({x_{j}})$, we have
@@
Example 3.18.
With the help of Example 3.12 and Theorem 3.17 we can bound the optimal redundancy of function $f$ defined in Example 3.9 as follows:
@@
3.2 Simplified Redundancy Lower Bounds
Using a smaller set of information vectors, one can obtain a lower bound on the redundancy of FCIDCs as follows:
Corollary 3.19.
Consider a collection of $M$ distinct binary words $x_{1},x_{2},\ldots,x_{M}$ of length $k$. Then, for a function $f$, the optimal redundancy of an FCIDCs is lower bounded as follows: